# GeoLangBind: Unifying Earth Observation with Agglomerative Vision–Language Foundation Models

### v1

@@ Appendix A Supplementary material
A.1 Details of the evaluation datasets
We evaluate GeoLangBind on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
The following datasets are part of the SkyScript [50] evaluation framework, which we follow for direct comparison.
  • SkyScript classification dataset [50]: contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
  • AID [54]: contains 2,000 aerial RGB images and 30 image scene classes.
  • EuroSAT [19]: contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
  • fMoW [11]: contains 106,081 RGB aerial images with 62 image scene classes.
  • Million-AID [30]: contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
  • PatternNet [68]: contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
  • NWPU-RESISC45 [10]: contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
  • RSI-CB [25]: contains 24,747 aerial RGB images and 35 image scene classes.
@@ A.1 Details of the evaluation datasets
Segmentation m-pv4ger-seg 320 x 320 2 3000 403 403 3 RGB
m-chesapeake-landcover 256 x 256 7 3000 1000 1000 4 RGBN
m-cashew-plantation 256 x 256 7 1350 400 50 13 Sentinel-2
m-SA-crop-type 256 x 256 10 3000 1000 1000 13 Sentinel-2
m-nz-cattle 500 x 500 2 524 66 65 3 RGB
@@
Additionally, we choose the GEO-Bench [24] suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations. Tab. 10 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
@@
In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) [23] dataset. The MADOS dataset is a high-resolution multispectral Sentinel-2 dataset for marine pollution detection. Spanning 174 globally distributed scenes (2015–2022) with 1.5M annotated pixels, it covers diverse pollutants, sea surface features, and water-related classes. Unlike existing datasets, MADOS enables scalable, generalizable deep learning models for holistic marine pollution monitoring.
A.2 Details on constructing GeoLangBind-2M
@@
A.2.1 Flair2-RGB-caption
For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet [63] dataset construction. We select the training set of 61,711 images from FLAIR #2 [15] and derive contextual information from their corresponding pixel-level segmentation masks. This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B [2] language model.
@@ A.2.1 Flair2-RGB-caption
A.2.2 Flair2-Elevation-caption
@@ A.2.3 MMflood-SAR-caption
A.2.4 NLCD-hyper-caption
@@
A.2.5 SAR-ship-caption and IR-ship-caption
For the SAR-ship-caption and IR-ship-caption subsets, we derive short captions from existing datasets to ensure high-quality textual descriptions. For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in EarthGPT [66], which is originally designed for visual question answering tasks. To generate descriptions for infrared images, we concatenate the corresponding question and answer pairs, transforming them into concise yet informative captions.
Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID [51], SSDD [65], and AIR-SARShip-2.0 [42]. These datasets provide extensive annotated SAR imagery, enabling the creation of meaningful captions. For further details on the captioning methodology, please refer to [66].
A.2.6 ChatEarthNet captions
@@
Regarding the ChatEarthNet-S2-caption and ChatEarthNet-SAR-caption datasets, we reuse the detailed descriptions provided in the original ChatEarthNet dataset. Specifically, we use Pixtral 12B [2] to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks. The word cloud for ChatEarthNet-caption datasets is presented in Fig. 14. To provide more insights into the dataset, we also show the histogram in Fig. 15. The prompt is as follows: {mdframed}[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
A.3 More dataset examples
@@
A.4 Ablation studies on distillation loss terms
@@
Tab. 11 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets. The baseline model (no distillation) achieves 59.10% (AID) and 32.96% (EuroSAT). Adding SigLIP distillation ($\alpha_{s}=1.0$) improves performance, while combining SigLIP and DinoV2 ($\alpha_{s}=1.0,\alpha_{d}=1.0$) further enhances results, especially for AID (65.36%). Introducing ViT distillation ($\alpha_{v}=1.0$) boosts EuroSAT accuracy, peaking at 38.63% when $\alpha_{s}=\alpha_{d}=\alpha_{v}=1.0$. In this work, we adopt the best overall configuration, $\alpha_{s}=2.0,\alpha_{d}=1.0,\alpha_{v}=1.0$, achieves the highest AID accuracy (69.20%) and strong EuroSAT performance (36.67%). This demonstrates the effectiveness of multi-teacher distillation.
A.5 Confusion matrices of zero-shot classification tasks
@@
Figs. 18, 19, 20, 21, 22, 24, 23 and 25 present the confusion matrices of our GeoLangBind-L-384 model on all eight datasets for zero-shot classification. These confusion matrices reveal that common model mistakes largely occur on classes for which humans would also have trouble distinguishing, or for which classes have similar semantic meaning. For instance, the model commonly confuses “sparse residential” and “medium residential” (RESISC45), “annual crop” and “permanent crop” (EuroSAT), “aquaculture land” and “pond” (SkyScript), “chaparral” and “desert” (RESISC45), “mine” and “quarry” (SkyScript, Million-AID), “recreational facility” and “stadium” and “playground” (fMoW, AID), and “ship” and “harbor” and “ferry terminal” (RESISC45, PatternNet).
@@
In Fig. 26, we show the comparison of confusion matrices between SkyScript (ViT-L) and GeoLangBind-L-384 on fine-grained zero-shot classification tasks. It can be clearly seen that our model performs better.
@@
Figure 18: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on SkyScript dataset.
@@
Figure 19: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on AID dataset.
@@
Figure 20: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on EuroSAT dataset.
@@
Figure 21: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on fMoW dataset.
@@
Figure 22: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on Million-AID dataset.
@@
Figure 23: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on PatternNet dataset.
@@
Figure 24: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on the RESISC45 dataset.
@@
Figure 25: Confusion matrix of the zero-shot classification results of GeoLangBind-L-384 on RSICB dataset.
@@
Figure 26: Comparison of confusion matrices between SkyScript-Large and GeoLangBind-L-384 on the fine-grained zero-shot classification task.
A.6 More visualization of features
@@
To better understand the representations learned by different models, Fig. 27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench. The m-chesapeake dataset primarily focuses on land cover classification, while the m-nz-cattle dataset captures cattle in remote sensing images. The feature visualization highlights key differences in spatial structure, consistency, and detail preservation across models.
@@
GeoLangBind produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript. The feature maps exhibit well-defined structures that align with meaningful spatial regions in the original images, such as roads, buildings, and vegetation.
@@
The feature maps from GeoLangBind demonstrate greater consistency across images with similar semantic content. As observed, buildings and roads consistently exhibit high activations, while tree-covered areas show lower activations. This structured activation pattern is apparent across multiple samples, suggesting that GeoLangBind learns more stable and semantically aligned representations.
@@
GeoLangBind retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines. This advantage is particularly evident in the m-nz-cattle dataset, where cattle are localized with sharper feature activations, as highlighted by the red bounding boxes. RemoteCLIP and SkyScript, on the other hand, exhibit more diffused and less distinct feature responses, making object-level distinctions less precise.
Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and GeoLangBind-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.

### v2

@@ Conclusion
Details of the evaluation datasets
We evaluate DOFA-CLIP on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
The following datasets are part of the SkyScript (Wang et al. 2024c) evaluation framework, which we follow for direct comparison.
  • SkyScript classification dataset (Wang et al. 2024c): contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
  • AID (Xia et al. 2017): contains 2,000 aerial RGB images and 30 image scene classes.
  • EuroSAT (Helber et al. 2019): contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
  • fMoW (Christie et al. 2018): contains 106,081 RGB aerial images with 62 image scene classes.
  • Million-AID (Long et al. 2021): contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
  • PatternNet (Zhou et al. 2018): contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
  • NWPU-RESISC45 (Cheng, Han, and Lu 2017): contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
  • RSI-CB (Li et al. 2017): contains 24,747 aerial RGB images and 35 image scene classes.
@@
Segmentation m-nz-cattle 500 x 500 2 524 66 65 3 RGB
@@
Additionally, we choose the GEO-Bench (Lacoste et al. 2024) suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations. Table 9 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
@@
In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) (Kikaki et al. 2024) dataset. The MADOS dataset is a high-resolution multispectral Sentinel-2 dataset for marine pollution detection. Spanning 174 globally distributed scenes (2015–2022) with 1.5M annotated pixels, it covers diverse pollutants, sea surface features, and water-related classes. Unlike existing datasets, MADOS enables scalable, generalizable deep learning models for holistic marine pollution monitoring.
Details on constructing GeoLangBind-2M
@@
Flair2-RGB-caption
For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet (Yuan et al. 2024) dataset construction. We select the training set of 61,711 images from FLAIR #2 (Garioud et al. 2023) and derive contextual information from their corresponding pixel-level segmentation masks. This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B (Agrawal et al. 2024) language model.
@@
Flair2-Elevation-caption
@@
NLCD-hyper-caption
@@
SAR-ship-caption and IR-ship-caption
For the SAR-ship-caption and IR-ship-caption subsets, we derive short captions from existing datasets to ensure high-quality textual descriptions. For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in Zhang et al. (2024a), which is originally designed for visual question answering tasks. To generate descriptions for infrared images, we concatenate the corresponding question and answer pairs, transforming them into concise yet informative captions.
Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID (Wei et al. 2020), SSDD (Zhang et al. 2021), and AIR-SARShip-2.0 (Sun, Sun, and Wang 2020). These datasets provide extensive annotated SAR imagery, enabling the creation of meaningful captions. For further details on the captioning methodology, please refer to Zhang et al. (2024a).
ChatEarthNet captions
@@
Regarding the ChatEarthNet-S2-caption and ChatEarthNet-SAR-caption datasets, we reuse the detailed descriptions provided in the original ChatEarthNet dataset. Specifically, we use Pixtral 12B (Agrawal et al. 2024) to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks. The word cloud for ChatEarthNet-caption datasets is presented in Fig. 14. To provide more insights into the dataset, we also show the histogram in Fig. 15. The prompt is as follows: {mdframed}[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
More dataset examples
@@
Ablation studies on distillation loss terms
@@
Table 10 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets. The baseline model (no distillation) achieves 59.10% (AID) and 32.96% (EuroSAT). Adding SigLIP distillation ($\alpha_{s}=1.0$) improves performance, while combining SigLIP and DinoV2 ($\alpha_{s}=1.0,\alpha_{d}=1.0$) further enhances results, especially for AID (65.36%). Introducing ViT distillation ($\alpha_{v}=1.0$) boosts EuroSAT accuracy, peaking at 38.63% when $\alpha_{s}=\alpha_{d}=\alpha_{v}=1.0$. In this work, we adopt the best overall configuration, $\alpha_{s}=2.0,\alpha_{d}=1.0,\alpha_{v}=1.0$, achieves the highest AID accuracy (69.20%) and strong EuroSAT performance (36.67%). This demonstrates the effectiveness of multi-teacher distillation.
Confusion matrices of zero-shot classification tasks
@@
Figure 18 to Fig. 25 present the confusion matrices of our DOFA-CLIP-L-384 model on all eight datasets for zero-shot classification. These confusion matrices reveal that common model mistakes largely occur for classes that humans would also have trouble distinguishing, or for classes with similar semantic meanings. For instance, the model commonly confuses “sparse residential” and “medium residential” (RESISC45), “annual crop” and “permanent crop” (EuroSAT), “aquaculture land” and “pond” (SkyScript), “chaparral” and “desert” (RESISC45), “mine” and “quarry” (SkyScript, Million-AID), “recreational facility” and “stadium” and “playground” (fMoW, AID), and “ship” and “harbor” and “ferry terminal” (RESISC45, PatternNet).
@@
In Fig. 26, we show the comparison of confusion matrices between SkyScript (ViT-L) and DOFA-CLIP-L-384 on fine-grained zero-shot classification tasks. It can be clearly seen that our model performs better.
@@
Figure 18: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on SkyScript dataset.
@@
Figure 19: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on AID dataset.
@@
Figure 20: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on EuroSAT dataset.
@@
Figure 21: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on fMoW dataset.
@@
Figure 22: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on Million-AID dataset.
@@
Figure 23: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on PatternNet dataset.
@@
Figure 24: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on the RESISC45 dataset.
@@
Figure 25: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on RSICB dataset.
@@
Figure 26: Comparison of confusion matrices between SkyScript-Large and DOFA-CLIP-L-384 on the fine-grained zero-shot classification task.
More visualization of features
@@
To better understand the representations learned by different models, Fig. 27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench. The m-chesapeake dataset primarily focuses on land cover classification, while the m-nz-cattle dataset captures cattle in remote sensing images. The feature visualization highlights key differences in spatial structure, consistency, and detail preservation across models.
@@
DOFA-CLIP produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript. The feature maps exhibit well-defined structures that align with meaningful spatial regions in the original images, such as roads, buildings, and vegetation.
@@
The feature maps from DOFA-CLIP demonstrate greater consistency across images with similar semantic content. As observed, buildings and roads consistently exhibit high activations, while tree-covered areas show lower activations. This structured activation pattern is apparent across multiple samples, suggesting that DOFA-CLIP learns more stable and semantically aligned representations.
@@
DOFA-CLIP retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines. This advantage is particularly evident in the m-nz-cattle dataset, where cattle are localized with sharper feature activations, as highlighted by the red bounding boxes. RemoteCLIP and SkyScript, on the other hand, exhibit more diffused and less distinct feature responses, making object-level distinctions less precise.
Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.